# Macroscopic entanglement between localized domain walls inside a cavity

### v1

@@ Abstract
We present a scheme for generating stable and tunable entanglement between two localized Bloch domain walls in nanomagnetic strips kept inside a chiral optical cavity. The entanglement is mediated by the effective optomechanical interaction between the cavity photons and the two macroscopic, collective modes of the pinned domain walls. By controlling the pinning potential and optical driving frequency, the robust, steady-state entanglement between the two macroscopic domain walls can survive beyond the typical milli-Kelvin temperature range.
@@ I Introduction
The field of quantum magnonics [1, 2] has seen rapid growth, driven by the quest to harness the coherent properties of quantized collective spin waves in magnetic materials or magnons for emerging quantum technologies and solid-state device applications. For instance, hybrid quantum systems [3, 4, 5] that combine magnons with other quantum systems such as photonic cavities or superconducting circuits provide a platform to perform a variety of quantum operations [6], ranging from quantum computing [7] and storage [8] to quantum transduction [9].
While much of the current focus in quantum magnonics has been on propagating magnons [10], localized magnetic spin textures such as chiral domain walls [11, 12], skyrmions [13], and spin vortices [14] can offer additional advantages for quantum computing and related applications. These textures can be spatially pinned [15, 16, 17] and arranged [18], readily observed [19, 20], and also transported in magnetic racetrack circuits [21]. Furthermore, they often exhibit rich topological properties [22, 23], making them attractive for both classical [24, 25, 26] and quantum [27, 28] information processing.
An essential task in implementing key quantum protocols is to generate and control entanglement, which has been extensively explored in propagating magnon modes [29, 30, 31, 32, 33, 34]. Most of these schemes rely on strong nonlinearity to generate entanglement either using Kerr effect [30], squeezed light [32] or optomechanical coupling in nonstationary magnon states [31]. However, generating stable, steady-state entanglement in localized spin textures remains largely unexplored. Recent experimental progress has significantly improved the ability to manipulate spin textures such as a Bloch domain wall (DW), especially in uniaxial magnetic materials [35]. These spin textures are already widely used in classical spintronic and nanoelectronic applications [36, 37] and can operate at much lower energies compared to their electronic counterparts. Importantly, in the quantum regime these DW behave as a macroscopic quasiparticle [38], and extending their control into the quantum domain, especially by inducing entanglement and squeezing simply through external driving, opens up exciting new opportunities for quantum information processing, computing, and memory applications.
@@
Figure 1: Experimental setup. The illustration shows a linearly polarized light created by controlling waveplates, which is passed through a Faraday isolator and quarter wave plate to excite circularly polarized optical mode. This creates a standing wave inside the cavity which is coupled to two ferromagnetic strips. The Bloch walls in the strips are pinned where the cavity field is present. The reflected optical mode is then collected at photodiode (PD) and analyzed via a spectrum analyzer. The output and input are separated using a beams-splitter (BS) and waveplate arrangement.
@@ II Setup and theoretical model
Magnetic spin textures such as a domain wall (DW) occur naturally in a range of magnetic materials [41, 42, 43] and have rich topological structure that depend on the bulk properties and can majorly be classified in two types: Bloch walls and Neel walls [44]. These DWs can be localized spatially or pinned using various mechanisms [45, 46, 47, 48]. Secondly, chiral cavities [49] are designed to excite a single polarization state such as left or right circularly polarized light. Such single-handed cavities have recently been demonstrated [39] and used for generating strong coupling with chiral molecules [50]. An illustration of the setup involving Bloch DWs in ferromagnetic strips inside an optical cavity is shown in Fig. 1. The magnetic strips are placed in such a way that the polarized light inside the cavity passes through the center of the pinned Bloch DWs. The radiation pressure forces the pinned magnon modes of the DW to oscillate due to the inverse Faraday effect. For a strong pinning potential, the magnon mode behaves like a collective macroscopic particle and can be canonically quantized to resemble a quantum harmonic oscillator [38].
@@
where $g_{j}$ is the coupling between a photon and the $j^{th}$ magnon mode arising due to inverse Faraday effect and $\xi$ is the strength of the driving field with frequency $\omega_{L}$. The magneto-optic coupling term is given by $g_{j}=-c\phi_{F}\sqrt{\epsilon}S^{j}_{\rm eff}/2$, and creates an effective optomechanical interaction between the cavity and magnon modes. This is independent of the cavity frequency $\omega_{a}$, unlike the usual optomechanical coupling where $g_{j}\propto\omega_{a}$ [40], and is dependent on the Faraday rotation $\phi_{F}$, which is the amount of rotation in polarization per unit length as light passes through a material of dielectric permittivity $\epsilon$. $S^{j}_{\rm eff}$ is a dimensionless geometrical factor that depends on the scattering cross-section, cavity volume and zero point fluctuation. See Appendix A for more details. The higher-order terms and non-linear Kerr effect have been neglected. In the frame moving with the driving field i.e., under the unitary transformation $\mathcal{H^{\prime}}=\mathcal{U}\mathcal{H}\mathcal{U}^{\dagger}-i\hbar\mathcal{U}\frac{\partial\mathcal{U}^{\dagger}}{\partial t}$, where $\mathcal{U}=e^{i\omega_{L}a^{\dagger}at}$, the Hamiltonian is time-independent
@@
Figure 2: Dissipative phases and two-mode entanglement of the system as a function of the detuning $\tilde{\Delta}$ and $G_{\rm{eff}}=G|g_{1}|/\omega_{1}$ for $\omega_{2}=\omega_{1}$ (top row) and $\omega_{2}=10~\omega_{1}$ (bottom row). The plots (a) and (e) show the number of stable solutions, with the distinct dissipative phases shown in (b) and (f). The entanglement between the two DWs $\mathcal{E}_{1|2}$ and between the photon and DW modes $\mathcal{E}_{a|1(2)}$ are shown in plots (c) and (g). Finally, the plots (d) and (h) show the entanglement between the photon and DW modes $\mathcal{E}_{a|1}$ and $\mathcal{E}_{a|2}$ for $\omega_{2}=10~\omega_{1}$. The blue and green dashed lines highlight two interesting regions of phase transition. The former indicates the jump from a region with two to three real solutions, while the latter exhibits a line of transcritical bifurcation where $\mathcal{E}_{1|2}$ in plot (c) is optimal. The parameters taken are $\kappa_{a}=2$ MHz, $\kappa_{1(2)}=1$ MHz, $\omega_{1}=1$ GHz, $g_{1(2)}=1$ MHz and temperature $T=2$ mK.
@@ III Stable Entangled Phases
Since our goal is to generate stable steady-state entanglement, it is important to first identify the driven dissipative phases of the system with stable solutions [51]. To this effect, the Hamiltonian in Eq. (3) is linearized around the steady state photon number $\bar{n}_{a}$. In the limit of strong driving, $\bar{n}_{a}$ is obtained as the roots of the cubic equation $\bar{n}_{a}=|\xi|^{2}/\left[\left(\Delta_{a}+\Omega\bar{n}_{\alpha}\right)^{2}+\kappa_{a}^{2}\right]$, where $\Omega=\sum_{j=1}^{2}2g^{2}_{j}\omega_{j}/(\omega^{2}_{j}+\kappa^{2}_{j})$, and $\kappa_{a}$ and $\kappa_{j}$ are the dissipation rates in cavity and $j^{\rm{th}}$ DW mode, respectively. The steady states of the different operators are written as $\langle a\rangle=\alpha^{(k)}$, $\langle a^{\dagger}a\rangle=\bar{n}_{a}^{(k)}$, and $\langle b_{j}\rangle=\beta_{j}^{(k)}$, where $k\in\{0,1,2\}$ indicate the roots (see Appendix B for the derivation). Let $G^{2}\equiv\bar{n}^{(k=0)}_{a}$ denote the real root common to all parameter regimes, with $G={|\xi|}/{\sqrt{\tilde{\Delta}^{2}+\kappa_{a}^{2}}}$, where $\tilde{\Delta}=\Delta_{a}+\Omega G^{2}$. The advantage of this representation is that $\bar{n}_{a}^{(k)}$ corresponding to the other roots can be parametrized in terms of $G$ and $\tilde{\Delta}$, such that for $k\in\{1,2\}$
@@
where $\tilde{\Delta}^{(k)}=\Delta_{a}-2\text{Re}(\beta_{j}^{(k)})$. To analyze the stability of the dissipative phases, we study the linearized quantum Langevin equations (QLE) [40] for $u=\left[\{x_{i},y_{i}\}\right]$ for $i\in\{a,1,2\}$, where $x_{i}=(\mathcal{O}_{i}+\mathcal{O}^{\dagger}_{i})/\sqrt{2}$ and $y_{i}=i(\mathcal{O}_{i}-\mathcal{O}^{\dagger}_{i})/\sqrt{2}$ are the quadrature operators for $\mathcal{O}_{i}\in[a,b_{1},b_{2}]$. For the $k^{th}$ root, the QLE is given by $\dot{u}^{(k)}=A^{(k)}u^{(k)}+\eta$, where $\eta=\left[\{\sqrt{2\kappa_{i}}~x^{\rm in}_{i},\sqrt{2\kappa_{i}}~p^{\rm in}_{i}\}\right]$ and
@@
Figure 2 illustrates the stable phases of the driven-dissipative system and the entanglement between the different modes, respectively. These phases are for different values of effective coupling $G_{\rm{eff}}=G|g_{1}|/\omega_{1}$ and the detuning $\tilde{\Delta}$, which can be controlled by the driving field strength and frequency. Moreover, we consider the cases where the two DW magnon modes are either resonant ($\omega_{2}=\omega_{1}$) or considerably detuned ($\omega_{2}=10\omega_{1}$), depending on the choice of the pinning field strength at the magnetic strips. The number of stable phases and the different dissipative phases are shown in Figs. 2(a,e) and (b,f), respectively, while Figs. 2(c) and (d,g,h) shows the DW-DW and photon-DW entanglement in the different phases of the system, maximized over all stable roots. Now, the key objective of the work is to demonstrate significant entanglement between the two DW magnon modes. In this regard, it is worth exploring regions where transcritical bifurcation occurs i.e., point where two of the roots exchange their stability (the stable root becomes unstable, while unstable root becomes stable), thus keeping the total number of stable roots fixed and marking a continuous dissipative phase transition. Let us consider the case $\bar{n}_{a}^{(0)}=\bar{n}_{a}^{(k)}$ for $k\in\{1,2\}$, where such a transition occurs at $G=\sqrt{-(\kappa_{a}^{2}+\tilde{\Delta}^{2})/2\tilde{\Delta}\Omega}$ for $\tilde{\Delta}\leq 0$. These bifurcations are highlighted in Fig. 2 by a green dashed line. The entanglement properties of the system along this line are remarkable. For $\omega_{1}=\omega_{2}$, the entanglement between photonic modes and the domain wall, $\mathcal{E}_{a|j}$ for $j\in\{1,2\}$, is maximum close to $\tilde{\Delta}=-\omega_{1}$, as shown in Figs. 2(c,d). Note that the photon mode is equally entangled with both the DW modes. However, moving away to the left along the transition line ($\tilde{\Delta}<0$), the photon-DW entanglement decreases and entanglement between the two DW modes $\mathcal{E}_{1|2}$ becomes significantly stronger. This transition from photon-DW to entanglement between the two magnetic textures is captured in Fig. 3(a). Incidentally, for unequal pinning strengths $\omega_{2}=10\omega_{1}$, the entanglement between the two DWs is suppressed. Moreover, the photon mode entangles with only one of the DW modes. For instance in Fig. 3(b), the photon mode entangles with the first DW mode at $\tilde{\Delta}=-\omega_{1}$ i.e., $\mathcal{E}_{a|1}$ is finite. Again by moving along the bifurcation line towards $\tilde{\Delta}=-\omega_{2}$, the entanglement can switch to the second DW mode and $\mathcal{E}_{a|2}$ is now significant. As such, by controlling the pinning strengths and the detuning, entanglement can be generated between the two DW modes or between the photon and a specific DW mode. Note that the tradeoff of entanglement is enforced by the monogamy principle [53, 54], which restricts the presence of significant entanglement simultaneously between all possible pairs of modes.

### v2

@@ Abstract
We present a scheme for generating stable and tunable entanglement between two localized Bloch domain walls in nanomagnetic strips kept inside a chiral optical cavity. The entanglement is mediated by the effective optomechanical interaction between the cavity photons and the two macroscopic, collective modes of the pinned domain walls. By control of the pinning potential and optical driving frequency, the robust, steady-state entanglement between the two macroscopic domain walls can survive beyond the typical millikelvin temperature range.
@@ I Introduction
The field of quantum magnonics [1, 2] has seen rapid growth, driven by the quest to harness the coherent properties of quantized collective spin waves in magnetic materials or magnons for emerging quantum technologies and solid-state device applications. For instance, hybrid quantum systems [3, 4, 5, 6, 7] that combine magnons with other quantum systems such as photonic cavities or superconducting qubits [8, 9, 10] provide a platform to perform a variety of quantum operations [11], ranging from quantum computing [12] and storage [13] to quantum transduction [14].
While much of the current focus in quantum magnonics has been on propagating magnons [15], localized magnetic spin textures such as chiral domain walls [16, 17], skyrmions [18, 19], and spin vortices [20] can offer additional advantages for quantum computing and related applications. These textures can be spatially pinned [21, 22, 23] and arranged [24], readily observed [25, 26], and also transported in magnetic racetrack circuits [27, 28]. Furthermore, they often exhibit rich topological properties [29, 30], making them attractive for both classical [31, 32, 33] and quantum [34, 35, 36] information processing.
An essential task in implementing key quantum protocols is to generate and control entanglement, which has been extensively explored in propagating magnon modes [37, 38, 39, 40, 41, 42]. Most of these schemes rely on strong nonlinearity to generate entanglement either using Kerr effect [38], squeezed light [40] or optomechanical coupling in nonstationary magnon states [39]. However, generating stable, steady-state entanglement in localized spin textures remains largely unexplored. Recent experimental progress has significantly improved the ability to manipulate spin textures such as a Bloch domain wall (DW), especially in uniaxial magnetic materials [43]. These spin textures are already widely used in classical spintronic and nanoelectronic applications [44, 45] and can operate at much lower energies compared to their electronic counterparts. Importantly, in the quantum regime, these DW behave as a macroscopic quasiparticle [46], and extending their control into the quantum domain, especially by inducing entanglement and squeezing simply through external driving, opens up exciting new opportunities for quantum information processing, computing, and memory applications.
@@
Figure 1: Experimental setup of domain walls inside a chiral cavity. The illustration shows linearly polarized light generated by a laser, which is passed through a Faraday isolator and a quarter-wave plate (QWP) to excite a circularly polarized optical mode. A right-circularly-polarized standing wave is created inside the chiral cavity, which then couples to the two ferromagnetic strips. The Bloch walls in the strips are pinned where the cavity field is present. The reflected optical mode is then collected at one port of a differential photodiode (Diff. PD) and the reference laser light is collected at the other port to obtain the spectrum via homodyne detection. The output and input are separated and merged with the use of a beam splitter (BS).
@@ II Setup and theoretical model
Magnetic spin textures such as a domain wall (DW) occur naturally in a range of magnetic materials [49, 50, 51] and have rich topological structure that depend on the bulk properties and can majorly be classified in two types: Bloch walls and Neel walls [52]. These DWs can be localized spatially or pinned using various mechanisms [53, 54, 55, 56]. Secondly, chiral cavities [57] are designed to excite a single polarization state such as left or right circularly polarized light. Such single-handed cavities have recently been demonstrated [47] and used for generating strong coupling with chiral molecules [58]. An illustration of the setup involving Bloch DWs in ferromagnetic strips inside an optical cavity is shown in Fig. 1. The magnetic strips are placed in such a way that the polarized light inside the cavity passes through the center of the pinned Bloch DWs. The magneto-optical interaction between the cavity and the domain wall arises due to radiation pressure, as the circularly polarized light in the cavity generates an effective magnetic field in the medium, which can then couple to the spins inside the DW. This is known as the inverse Faraday effect [59]. While spin-orbit coupling mediates such effect in electronic systems [60], the interaction between the DW and cavity modes can be better understood in terms of how the electric permittivity tensor is modulated by spin oscillations. For a strong pinning potential, the DW mode behaves like a collective macroscopic particle and can be canonically quantized to resemble a quantum harmonic oscillator [46].
Figure 2: Dissipative phases and two-mode entanglement of the system as a function of the detuning $\tilde{\Delta}$ and $G_{\rm{eff}}=G|g_{1}|/\omega_{1}$ for $\omega_{2}=\omega_{1}$ (a-d) and $\omega_{2}=10~\omega_{1}$ (e-h). The plots (a) and (e) show the number of stable solutions, with the distinct dissipative phases $p\in\{0,1,...,6\}$ shown in (b) and (f). The entanglement between the two DWs $\mathcal{E}_{1|2}$ and between the photon and DW modes $\mathcal{E}_{a|1(2)}$ are shown in plots (c) and (d). Finally, the plots (g) and (h) show the entanglement between the photon and DW modes $\mathcal{E}_{a|1}$ and $\mathcal{E}_{a|2}$ for $\omega_{2}=10~\omega_{1}$. The blue and green dashed lines highlight two interesting regions of phase transition. The former indicates the jump from a region with two to three real solutions, while the latter exhibits a line of transcritical bifurcation where $\mathcal{E}_{1|2}$ in plot (c) is optimal. The parameters taken are $\kappa_{a}=2$ MHz, $\kappa_{1(2)}=1$ MHz, $\omega_{1}=1$ GHz, $g_{1(2)}=1$ MHz and temperature $T=2$ mK.
@@
where $g_{j}$ is the coupling between a photon and the $j^{th}$ magnon mode arising due to inverse Faraday effect and $\xi$ is the strength of the driving field with frequency $\omega_{L}$. The magneto-optic coupling term is given by $g_{j}=-c\phi_{F}\sqrt{\epsilon}S^{j}_{\rm eff}/2$, and creates an effective optomechanical interaction between the cavity and magnon modes. This is independent of the cavity frequency $\omega_{a}$, unlike the usual optomechanical coupling where $g_{j}\propto\omega_{a}$ [48], and is dependent on the Faraday rotation $\phi_{F}$, which is the amount of rotation in polarization per unit length as light passes through a material of dielectric permittivity $\epsilon$. $S^{j}_{\rm eff}$ is a dimensionless geometrical factor that depends on the scattering cross-section, cavity volume and zero point fluctuation. See Appendices A and B for more details. The higher-order terms and non-linear Kerr effect have been neglected. In the frame moving with the driving field i.e., under the unitary transformation $\mathcal{H^{\prime}}=\mathcal{U}\mathcal{H}\mathcal{U}^{\dagger}-i\hbar\mathcal{U}\frac{\partial\mathcal{U}^{\dagger}}{\partial t}$, where $\mathcal{U}=e^{i\omega_{L}a^{\dagger}at}$, the Hamiltonian is time-independent
@@ III Stable Entangled Phases
Since our goal is to generate stable steady-state entanglement, it is important to first identify the driven dissipative phases of the system with stable solutions [61]. To this effect, the Hamiltonian in Eq. (3) is linearized around the steady state photon number $\bar{n}_{a}$. In the limit of strong driving, $\bar{n}_{a}$ is obtained as the roots of the cubic equation $\bar{n}_{a}=|\xi|^{2}/\left[\left(\Delta_{a}+\Omega\bar{n}_{\alpha}\right)^{2}+\kappa_{a}^{2}\right]$, where $\Omega=\sum_{j=1}^{2}2g^{2}_{j}\omega_{j}/(\omega^{2}_{j}+\kappa^{2}_{j})$, and $\kappa_{a}$ and $\kappa_{j}$ are the dissipation rates in cavity and $j^{\rm{th}}$ DW mode, respectively. The steady states of the different operators are written as $\langle a\rangle=\alpha^{(k)}$, $\langle a^{\dagger}a\rangle=\bar{n}_{a}^{(k)}$, and $\langle b_{j}\rangle=\beta_{j}^{(k)}$, where $k\in\{0,1,2\}$ indicate the roots (see Appendix C for the derivation). Let $G^{2}\equiv\bar{n}^{(k=0)}_{a}$ denote the real root common to all parameter regimes, with $G={|\xi|}/{\sqrt{\tilde{\Delta}^{2}+\kappa_{a}^{2}}}$, where $\tilde{\Delta}=\Delta_{a}+\Omega G^{2}$. The advantage of this representation is that $\bar{n}_{a}^{(k)}$ corresponding to the other roots can be parametrized in terms of $G$ and $\tilde{\Delta}$, such that for $k\in\{1,2\}$
@@
where $\tilde{\Delta}^{(k)}=\Delta_{a}-2\text{Re}(\beta_{j}^{(k)})$. To analyze the stability of the dissipative phases, we study the linearized quantum Langevin equations (QLE) [48] for $u=\left[\{x_{i},y_{i}\}\right]$ for $i\in\{a,1,2\}$, where $x_{i}=(\mathcal{O}_{i}+\mathcal{O}^{\dagger}_{i})/\sqrt{2}$ and $y_{i}=i(\mathcal{O}_{i}-\mathcal{O}^{\dagger}_{i})/\sqrt{2}$ are the quadrature operators for $\mathcal{O}_{i}\in[a,b_{1},b_{2}]$. See Appendix D for the equations.
For the $k^{th}$ root, the QLE is given by $\dot{u}^{(k)}=A^{(k)}u^{(k)}+\eta$, where $\eta=\left[\{\sqrt{2\kappa_{i}}~x^{\rm in}_{i},\sqrt{2\kappa_{i}}~p^{\rm in}_{i}\}\right]$ and
@@
Figure 2 illustrates the stable phases of the driven-dissipative system and the entanglement between the different modes, respectively. These phases are for different values of effective coupling $G_{\rm{eff}}=G|g_{1}|/\omega_{1}$ and the detuning $\tilde{\Delta}$, which can be controlled by the driving field strength and frequency, and obtained by solving the mean-field QLE. Moreover, we consider the cases where the two DW magnon modes are either resonant ($\omega_{2}=\omega_{1}$) or considerably detuned ($\omega_{2}=10\omega_{1}$), depending on the choice of the pinning field strength at the magnetic strips. The number of stable phases and the different dissipative phases are shown in Figs. 2(a,e) and (b,f), respectively, while Figs. 2(c) and (d,g,h) shows the DW-DW and photon-DW entanglement in the different phases of the system, maximized over all stable roots. The different phases in Fig. 2(b,f) are based on which of the three roots ($k\in\{0,1,2\}$) are stable. Using binary digits, we can identify whether the $k^{th}$ root is stable ($s_{k}=1$) or not ($s_{k}=0$). The dissipative phase is then given by the number $p=\sum_{k}2^{k}s_{k}$. As there are no phases where all three roots are stable, so the phases range from 0 up to 6. Note that the mean photon number $\bar{n}_{a}^{(k)}$ is an order parameter for the dissipative phase transitions [61].
@@
Now, the key objective of the work is to demonstrate significant entanglement between the two DW magnon modes. In this regard, it is worth exploring regions where transcritical bifurcation occurs i.e., point where two of the roots exchange their stability (the stable root becomes unstable, while unstable root becomes stable), thus keeping the total number of stable roots fixed and marking a continuous dissipative phase transition. Let us consider the case $\bar{n}_{a}^{(0)}=\bar{n}_{a}^{(k)}$ for $k\in\{1,2\}$, where such a transition occurs at $G=\sqrt{-(\kappa_{a}^{2}+\tilde{\Delta}^{2})/2\tilde{\Delta}\Omega}$ for $\tilde{\Delta}\leq 0$. These bifurcations are highlighted in Fig. 2 by a green dashed line. The entanglement properties of the system along this line are remarkable. For $\omega_{1}=\omega_{2}$, the entanglement between photonic modes and the domain wall, $\mathcal{E}_{a|j}$ for $j\in\{1,2\}$, is maximum close to $\tilde{\Delta}=-\omega_{1}$, as shown in Figs. 2(c,d). Note that the photon mode is equally entangled with both the DW modes. However, moving away to the left along the transition line ($\tilde{\Delta}<0$), the photon-DW entanglement decreases and entanglement between the two DW modes $\mathcal{E}_{1|2}$ becomes significantly stronger. This transition from photon-DW to entanglement between the two magnetic textures occurs at $\tilde{\Delta}/\omega_{1}=\kappa_{a}/\kappa_{1}$, and is captured in Fig. 3(a). Incidentally, for unequal pinning strengths $\omega_{2}=10\omega_{1}$, the entanglement between the two DWs is suppressed. Moreover, the photon mode entangles with only one of the DW modes. For instance in Fig. 3(b), the photon mode entangles with the first DW mode at $\tilde{\Delta}=-\omega_{1}$ i.e., $\mathcal{E}_{a|1}$ is finite. Again by moving along the bifurcation line towards $\tilde{\Delta}=-\omega_{2}$, the entanglement can switch to the second DW mode and $\mathcal{E}_{a|2}$ is now significant. As such, by controlling the pinning strengths and the detuning, entanglement can be generated between the two DW modes or between the photon and a specific DW mode. Note that the tradeoff of entanglement is enforced by the monogamy principle [63, 64], which restricts the presence of significant entanglement simultaneously between all possible pairs of modes.
A physical interpretation of the existence of entanglement along the bifurcation line can be provided by analyzing the correlations between the optical and DW modes i.e., the covariances $\langle x_{i}x_{j}\rangle$, $\langle x_{i}y_{j}\rangle$ and $\langle y_{i}y_{j}\rangle$ for $i\in\{a,1,2\}$. In the vicinity of dissipative phase transition, correlations and decay rates tend to diverge [65]. Our numerical results  [^1] All simulations and illustrations were done with the help of open source Python libraries [95] and repositories for the software Blender [96, 97], respectively. show that correlation between the cavity and DW modes $\langle p_{a}x_{1}\rangle$ or the DW-DW modes $\langle x_{1}x_{2}\rangle$ are often large and comparable to local mode fluctuations such $\langle p_{a}p_{a}\rangle$ or $\langle x_{1}x_{1}\rangle$. This allows for build up large quadrature correlations and entanglement, as quantified by the logarithmic negativity. A qualitative discussion on the variation of correlation and local fluctuation near the bifurcation line is presented in Appendix E.